OpenStudy (anonymous):

∫tan(sin^-1x)dx/ √ 1-x^2 let,tan(sin^-1x)=z or, d/dx tan(sin^-1x)=dz/dx or,sec^2(sin^-1x).1/√ 1-x^2=dz/dx or, dx/√ 1-x^2=dz/1-tan^2(sin^-1x) or,dx/√ 1-x^2=dz/1+z^2 ∫tan(sin^-1x)dx/ √ 1-x^2=∫z.dz/1+z^2=1/1+z^2∫zdz=1/1+z^2.z^1+1/1+1+c=1/1+z^2.z^2/2+c=1/1+(tansin^-1x)2.(tansin^-1x)^2/2+c=tan(sin^-1x)^2/2(1+(tansin^-1x)^2+c is this is right?

OpenStudy (austinl):

Okay first off, wow. That is a mess. You need to learn \(\LaTeX\) It will make everything much easier to understand. Take your first line for example. \(\Large{\int \frac{\tan{(\sin^{-1}{(x)})} dx}{\sqrt{1-x^2}}}\)

OpenStudy (anonymous):

\(u=\sin^{-1}(x)\) gets it

OpenStudy (austinl):

That is what I was thinking, a simple substitution. Very nice.

OpenStudy (anonymous):

long as your recall \[\int \tan(u)du=\ln(\sec(u))\] you are done in one step

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@Rohitkhanna